# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 3.6 Rigorous PP analysis
$\bar{\boldsymbol{U}}_{ij}^{(\mu)}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\geq\frac{-\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\left(\Delta y\left(B_{1}\right)_{i+\frac{1}{2},j}^{-,\mu}-\Delta y\left(B_{1}\right)_{i-\frac{1}{2},j}^{+,\mu}+\Delta x\left(B_{2}\right)_{i,j+\frac{1}{2}}^{\mu,-}-\Delta x\left(B_{2}\right)_{i,j-\frac{1}{2}}^{\mu,+}\right)}{\Delta y\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)+\Delta x\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)}.$
@@
$\Pi_{2}$ $\geq-\frac{\Delta t_{n}}{\Delta x\Delta y}\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\Delta y\left(B_{1}\right)_{i+\frac{1}{2},j}^{-,\mu}-\Delta y\left(B_{1}\right)_{i-\frac{1}{2},j}^{+,\mu}+\Delta x\left(B_{2}\right)_{i,j+\frac{1}{2}}^{\mu,-}-\Delta x\left(B_{2}\right)_{i,j-\frac{1}{2}}^{\mu,+}\right)$ (66)
$=-\Delta t_{n}\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)\text{div}_{\text{within}}\widetilde{\boldsymbol{B}}_{\sigma}^{n}=0,$
where the last equality uses the LDF property of the magnetic field. By using the inequality (60) in Corollary 2, we give a lower bound of $\Pi_{3}$ as
@@
$\Pi_{4}$ $\geq\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right)$ (68)
$+\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(-\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right)$
$+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right)$
$+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(-\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right).$
@@
$\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}$
$\geq\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{1}^{+}+\frac{\Delta t_{n}}{\Delta x}\left(\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}-\frac{\left|\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}}\right)\right)\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)$
$+\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{1}^{-}-\frac{\Delta t_{n}}{\Delta x}\left(\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}}\right)\right)\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)$
$+\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{2}^{+}+\frac{\Delta t_{n}}{\Delta y}\left(\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}-\frac{\left|\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}}\right)\right)\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)$
$+\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{2}^{-}-\frac{\Delta t_{n}}{\Delta y}\left(\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}\right)\right)\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)$
@@
where the last inequality uses the CFL condition (61) and PP limiting condition (54). Therefore, we have
$\bar{\rho}_{ij}^{n+1}>0,\quad\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}>0,\forall\boldsymbol{n}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3},$
@@
(PP via Zhang-Shu convex decomposition). Let the wave speeds in the HLL flux satisfy the condition (57). If the polynomial vectors $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ are locally divergence-free and satisfy the following PP limiting condition,
@@
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G},\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}^{\text{Zhang-Shu}},$
@@
where $\mathbb{S}_{ij}^{\text{Zhang-Shu}}$ denotes the set of all nodes of Zhang-Shu convex decomposition (48), then the scheme (55) preserves $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ under the CFL-type condition
$\Delta t_{n}\leq c_{0}\min\left\{\frac{a_{1}\Delta x\Delta y\hat{\omega}_{1}^{GL}}{\left(a_{2}\Delta x+a_{1}\Delta y\right)\max\left\{\alpha_{1,i+\frac{1}{2}}^{\mu},\alpha_{1,i-\frac{1}{2}}^{\mu}\right\}},\frac{a_{2}\Delta x\Delta y\hat{\omega}_{1}^{GL}}{\left(a_{2}\Delta x+a_{1}\Delta y\right)\max\left\{\alpha_{2,j+\frac{1}{2}}^{\mu},\alpha_{2,j-\frac{1}{2}}^{\mu}\right\}}\right\},$
$\quad\forall I_{ij}\in\Omega,1\leq\mu\leq q,$
@@
where $\hat{\omega}_{1}^{GL}=\frac{1}{L\left(L-1\right)}$ with $L=\lceil\frac{k+3}{2}\rceil$.
Corollary 4.
(PP via optimal convex decomposition). Let the wave speeds in the HLL flux satisfy the condition (57). If the polynomial vectors $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ are locally divergence-free and satisfy the following PP limiting condition,
@@
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G},\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}^{\text{optimal}},$
@@
where $\mathbb{S}_{ij}^{\text{optimal}}$ denotes the set of all nodes of optimal convex decomposition (49), then the scheme (55) preserves $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ under the CFL-type condition
$\Delta t_{n}\leq c_{0}\min\left\{\frac{a_{1}\Delta x\Delta y}{\max\left\{6a_{1}\Delta y+2a_{2}\Delta x,6a_{2}\Delta x+2a_{1}\Delta y\right\}\max\left\{\alpha_{1,i+\frac{1}{2}}^{\mu},\alpha_{1,i-\frac{1}{2}}^{\mu}\right\}},\right.$

### v2

@@ 3.6 Rigorous PP analysis
$+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)\left(\bar{\boldsymbol{U}}_{2,ij}^{(\mu)}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right),$
$\Pi_{3}:=$ $-\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}\left(\boldsymbol{H}_{1,i+\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)-\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}\left(\boldsymbol{H}_{1,i-\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right]$
$-\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}\left(\boldsymbol{H}_{2,j+\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)-\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}\left(\boldsymbol{H}_{2,j-\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right],$
$\Pi_{4}:=$ $-\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)\cdot\boldsymbol{n}^{*}+\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\cdot\boldsymbol{n}^{*}\right)$
$-\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)\cdot\boldsymbol{n}^{*}+\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\cdot\boldsymbol{n}^{*}\right).$
We now estimate the lower bounds of $\Pi_{0}$, $\Pi_{2}$, $\Pi_{3}$, and $\Pi_{4}$, respectively. Based on the the convex decomposition (50), we have
$\Pi_{0}=$ $\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\omega_{1}^{+}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)+\omega_{1}^{-}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)+\omega_{2}^{+}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right.$ (69)
$\left.+\omega_{2}^{-}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right]+\sum_{s=1}^{S}\omega_{s}\left(\widetilde{\boldsymbol{U}}_{\sigma}^{n}\left(x_{ij}^{(s)},y_{ij}^{(s)}\right)\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right),$
$\geq\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\omega_{1}^{+}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)+\omega_{1}^{-}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right.$
$\quad\left.+\omega_{2}^{+}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)+\omega_{2}^{-}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right],$
where the inequality follows from the condition (58) and Lemma 1. According to Lemma 3, for any $\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}$, we have
$\bar{\boldsymbol{U}}_{1,ij}^{(\mu)}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\geq\frac{-\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}}{\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}}\left(\left(B_{1}\right)_{i+\frac{1}{2},j}^{-,\mu}-\left(B_{1}\right)_{i-\frac{1}{2},j}^{+,\mu}\right),$
$\bar{\boldsymbol{U}}_{2,ij}^{(\mu)}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\geq\frac{-\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}}{\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}}\left(\left(B_{2}\right)_{i,j+\frac{1}{2}}^{\mu,-}-\left(B_{2}\right)_{i,j-\frac{1}{2}}^{\mu,+}\right).$
@@
$\Pi_{2}$ $\geq-\frac{\Delta t_{n}}{\Delta x}\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\left(B_{1}\right)_{i+\frac{1}{2},j}^{-,\mu}-\left(B_{1}\right)_{i-\frac{1}{2},j}^{+,\mu}\right)$ (70)
$\quad-\frac{\Delta t_{n}}{\Delta y}\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\left(B_{2}\right)_{i,j+\frac{1}{2}}^{\mu,-}-\left(B_{2}\right)_{i,j-\frac{1}{2}}^{\mu,+}\right)$
$=-\Delta t_{n}\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)\text{div}_{\text{local}}\widetilde{\boldsymbol{B}}_{\sigma}^{n}=0,$
where the last equality uses the LDF property of the magnetic field within $I_{ij}$:
$\text{div}_{\text{local}}\widetilde{\boldsymbol{B}}_{\sigma}^{n}$ $=\frac{1}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i+\frac{1}{2}}^{-},y_{j}^{(\mu)}\right)-{\left(B_{1}\right)}_{\sigma}^{n}\left(x_{i-\frac{1}{2}}^{+},y_{j}^{(\mu)}\right)\right]$
$\quad+\frac{1}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[{\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j+\frac{1}{2}}^{-}\right)-{\left(B_{2}\right)}_{\sigma}^{n}\left(x_{i}^{(\mu)},y_{j-\frac{1}{2}}^{+}\right)\right]$
$=\frac{1}{\Delta x\Delta y}\oint_{\partial I_{ij}}\widetilde{\boldsymbol{B}}_{ij}^{n}\left(\boldsymbol{x}\right)\cdot\boldsymbol{n}\mathrm{~{}d}s=\frac{1}{\Delta x\Delta y}\int_{I_{ij}}\nabla\cdot\widetilde{\boldsymbol{B}}_{ij}^{n}\left(\boldsymbol{x}\right)\mathrm{~{}d}\boldsymbol{x}=0.$
By using the inequality (65) in Lemma 4, we give a lower bound of $\Pi_{3}$ as
@@
$\Pi_{4}$ $\geq\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right)$
$+\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right)$
$+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right)$
$+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\left(\boldsymbol{v}^{*}\cdot\boldsymbol{B}^{*}\right)-\frac{\left|\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right).$ (72)
@@
$\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}$
$\geq\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{1}^{+}+\frac{\Delta t_{n}}{\Delta x}\left(\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}-\frac{\left|\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}}\right)\right)\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)$
$+\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{1}^{-}-\frac{\Delta t_{n}}{\Delta x}\left(\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}}\right)\right)\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)$
$+\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{2}^{+}+\frac{\Delta t_{n}}{\Delta y}\left(\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}-\frac{\left|\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}}\right)\right)\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)$
$+\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\omega_{2}^{-}-\frac{\Delta t_{n}}{\Delta y}\left(\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}\right)\right)\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)$
@@
where the last inequality follows from the condition (58) and CFL condition (66). In summary, we have
$\bar{\rho}_{ij}^{n+1}>0,\quad\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}>0\quad\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3},$
@@
Let the OEDG solution polynomial vectors $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ be LDF and satisfy the following condition after the PP limiting procedure:
@@
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G}\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}^{\text{classic}},$
@@
where $\mathbb{S}_{ij}^{\text{classic}}$ denotes the set of all nodes of the classic Zhang–Shu convex decomposition (51). If the wave speeds in the HLL flux satisfy the condition (62), then the scheme (59) preserves $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ under the CFL-type condition
$\Delta t_{n}\leq\frac{\Delta x\Delta y\hat{\omega}_{1}^{GL}}{a_{2}\Delta x+a_{1}\Delta y},$
@@
where $a_{1}=\max\{\alpha_{1,i+\frac{1}{2}}^{\mu}\}$, $a_{2}=\max\{\alpha_{2,j+\frac{1}{2}}^{\mu}\}$, and $\hat{\omega}_{1}^{GL}=\frac{1}{L\left(L-1\right)}$ with $L=\lceil\frac{k+3}{2}\rceil$.
Proof This follows from Theorem 1 by taking the convex decomposition as the Zhang–Shu convex decomposition (51) with $a_{1}=\max\{\alpha_{1,i+\frac{1}{2}}^{\mu}\}$ and $a_{2}=\max\{\alpha_{2,j+\frac{1}{2}}^{\mu}\}$. ∎
Corollary 2 (PP via optimal convex decomposition for $\mathbb{P}^{2}$ and $\mathbb{P}^{3}$).
Consider the $\mathbb{P}^{2}$- and $\mathbb{P}^{3}$-based LDF OEDG method. Let the OEDG solution polynomial vectors $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ be LDF and satisfy the following condition after the PP limiting procedure:
@@
$\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\in\mathcal{G}\quad\forall\boldsymbol{x}\in\mathbb{S}_{ij}^{\text{optimal}},$
@@
where $\mathbb{S}_{ij}^{\text{optimal}}$ denotes the set of all nodes of optimal convex decomposition (52). If the wave speeds in the HLL flux satisfy the condition (62), then the scheme (59) preserves $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ under the CFL-type condition
$\Delta t_{n}\leq\frac{\Delta x\Delta y}{\max\left\{6a_{1}\Delta y+2a_{2}\Delta x,6a_{2}\Delta x+2a_{1}\Delta y\right\}},$
where $a_{1}=\max\{\alpha_{1,i+\frac{1}{2}}^{\mu}\}$ and $a_{2}=\max\{\alpha_{2,j+\frac{1}{2}}^{\mu}\}$.
Proof This follows from Theorem 1 by adopting the optimal convex decomposition (52) with $a_{1}=\max\{\alpha_{1,i+\frac{1}{2}}^{\mu}\}$ and $a_{2}=\max\{\alpha_{2,j+\frac{1}{2}}^{\mu}\}$. ∎
Remark 3.
For the $\mathbb{P}^{k}$-based LDF OEDG method with higher $k\geq 4$, the PP CFL condition derived from the optimal convex decomposition can be expressed as
$\Delta t_{n}\left(\frac{a_{1}}{\Delta x}+\frac{a_{2}}{\Delta y}\right)\leq\bar{\omega}_{\star}(\delta,\mathbb{P}^{k}),$
where weight $\bar{\omega}_{\star}=\bar{\omega}_{\star}(\delta,\mathbb{P}^{k})$ depends on $\delta:=\frac{a_{1}\Delta y-a_{2}\Delta x}{a_{1}\Delta y+a_{2}\Delta x}$ and the space $\mathbb{P}^{k}$. For example, $\bar{\omega}_{\star}(\delta,\mathbb{P}^{1})=\frac{1}{2}$, $\bar{\omega}_{\star}(\delta,\mathbb{P}^{2})=\bar{\omega}_{\star}(\delta,\mathbb{P}^{3})=\frac{1}{4+2|\delta|}$, and for $k\in\{4,5\}$,